OpenStudy (anonymous):

A mirror with a parabolic cross section is used to collect sunlight on a pipe located at the focus of the mirror. The pipe is located 2 inches from the vertex of the mirror. Write an equation of the parabola that models the cross section of the mirror. Assume that the parabola opens upward.

OpenStudy (campbell_st):

so the focal length is 2 then the parabola is in the form \[y = \frac{1}{8}(x -h)^2 + k\] or in the form \[(x -h)^2 = 4 \times 2(y - k)\] with a vertex at (h, k)